IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

  • finished adding ubercal info to dvo
  • creating a test suite using the dvo shell test framework for the new zero it and relphot features
  • added some features to dvo shell to support the test frame work (better generic fits table output for vectors)
  • ran more test exposures for doug to check on the faint end bias (grizy for bad and good nights in MD04 & MD09)

Serge Chastel

  • Mon: holiday;
  • Tue & Wed: sick;
  • Nebulous replication: Changed neb-admin retarget and balance mysql statements to divide execution time by 10.

Heather Flewelling

  • vacation monday through wednesday
  • merging ThreePi.v3- seems to be working well and merging faster due to the changes (merged 26, 33 are left)
  • stacks for jt
  • various post conference/travel tasks

Roy Henderson

  • PSVO:
    • reinstated ecliptic and galactic plane plotting
    • finalized fixes for changing user and not breaking Derby connection
    • added nice auto-coloring to regex constrained textfields if users inputs wrong values
  • Support:
    • Robert Senger's email regarding multiple detections per object in the stack
    • questions from Tomo. Worked on a query for him to get apMags etc
    • more Robert Senger support emails (ippDetectIDs, nDetections question)
  • ippToPsps:
    • discussion with Gene regarding new way to load data from DVO
    • beginnings of a design for the above
    • looked into necessary gpc1 Db changes to access data in DVO
  • Documentation:
    • added lots of information to the PSPS FAQ wiki page

Mark Huber

  • caught up with DRAV group minutes
  • ppSub convolution direction testing, continuing to collect test samples. also
    • looking in detail at flagged events from QUB, found recent one not really a convolution direction issue
    • setting up tests for monitoring nightly SSdiffs
  • MD.GR0
    • MD06.V3 and MD04.V3 full reprocessing of nightly stacks finished and distributed to datastore
    • investigated small sample of edge skycell faults that had strangely large/small PSF
  • TSS event/name server test setup ipp0032 for consortium access

Bill Sweeney

  • Spent most of this week working on detectability queries. There are a number of problems related to update processing that must be resolved.
  • MPG found some M31 data that they did not receive. In the starting and stopping with the ipp064 data loss one of us got confused. Worked on regenerating that data.
  • debugged a ppImage crash that turned out to be due to bad data - an exposure where the dome was closed.
  • debugged psphotStack failures. Found problem in photometry that is now on Gene's list to investigate.

Chris Waters

  • Reallocated hosts in pantasks to better utilize the machines now that we no longer need to spend time magicking and destreaking data.
  • Kicked LAP repeatedly to get processing moving after bottlenecks (largely a result of ipp064 data loss).
  • Worked on making the new PATTERN.CONTINUITY correction functional in ppImage. This aims to ensure that the boundaries of adjacent cells have similar values, which will prevent problems from arising due to the background subtraction.